# AI Failure Loops in Devalued Work: The Confluence of Overconfidence in AI & Underconfidence in Worker Expertise

### v1

@@ 7 Discussion: Toward Positive Futures for AI and Work
The measurements used to evaluate AI models are but one of many representations of the world (Boyd and Crawford, 2012). Our case studies demonstrate how the design of workplace AI evaluation metrics can further marginalize devalued workers. As we described in Section 6.1, managers measured worker versus AI performance according to AI developers’ flawed predictive target. Even though this predictive target misaligned with workers’ underlying decision-making goals (Cheng et al., 2022), this comparative analysis was used to justify the AI tools’ deployment.
We should also question the value of “measuring” workers in the first place. As noted by CSCW scholars, the question of what work to make (in)visible through datafication can be politically fraught (Star and Strauss, 1999). Researchers and developers may find that some work cannot or should not be quantifiable, leading to the development of alternative evaluation standards (e.g., worker satisfaction). Even in occupations where there has been decades of research and development, as is the case with AIED technologies in K-12 education, the vast majority of evaluations for AI deployments have focused on understanding impacts on other stakeholders (e.g., students), without consideration for frontline workers (e.g., teachers’) experiences and perspectives (Holmes, 2023).
@@ 7.4 The Case of Feminized Labor
Redirecting toward a pro-worker future for AI may involve some unique considerations in the context of feminized labor. For instance, many feminized occupations center on emotional and care-oriented labor that may be difficult or impossible to quantify, increasing risks of misquantifications in AI development. Moreover, workers in service-oriented positions (i.e., most feminized positions) are often trained to decenter themselves, becoming a “non-person” in their workplace and to employers (Forsythe, 1993; Ming et al., 2023). Even if AI developers manage to form genuine design partnerships with workers, they may not—without explicit encouragement—advocate for designing technologies that meet their own needs. In past efforts to design technologies with feminized workers (e.g., home health aides, K-12 teachers), these workers often centered clients’ needs and de-centered their own needs by default during design engagements, unless nudged by designers to do so (holstein2019; Bartle et al., 2022). We anticipate that, with time, redirecting this trajectory will only increase in importance for AI in feminized occupations. When one considers that many of the most devalued occupations are dominated by women, the majority of whom are women of color, it becomes necessary to recognize the presence and prevalence of AI Failure Loops as a racial, gender, and social justice issue.
@@ 8 Conclusion
Visions for “transformative” workplace AI technologies that undermine workers are not new. Researchers and practitioners have long hoped AI technologies will “fix” staff shortages in feminized occupations. The challenges described in this paper—the development of workplace technology without a grounded understanding of workers—are reminiscent of observations made by Forsythe nearly three decades ago (Forsythe, 1993). Forsythe found that AI developers making expert systems held impoverished understandings of “work,” disregarding social and maintenance activities. Thus, these expert systems failed to model experts’ real-world work practices. Despite the shift from rule-based to machine learning systems, many of Forsythe’s observations from three decades ago persist today.
@@
We would like to thank the reviewers for carefully guiding us in strengthening our submission. We are also grateful for the thoughtful feedback shared by Alicia Devrio, Jodi Forlizzi, Nari Johnson, Joy Ming, Tina Park, and David Widder.

### v2

@@ 7 Discussion: Toward Pro-Worker AI
The measurements used to evaluate AI models are but one of many representations of the world (Boyd and Crawford, 2012). Our case studies demonstrate how the design of workplace AI evaluation metrics can further marginalize devalued workers. As we described in Section 6.1, managers measured worker versus AI performance according to AI developers’ flawed predictive target. Even though this predictive target misaligned with workers’ underlying decision-making goals (Cheng et al., 2022), this comparative analysis was used to justify the AI tools’ deployment. Moreover, once AI tools were deployed, these flawed measurements were used to track and surveil workers’ compliance with AI outputs (Kawakami et al., 2022a).
Our case studies also strongly suggest that we should question the value of “measuring” workers in the first place. As noted by CSCW scholars, the question of what work to make (in)visible through datafication can be politically fraught (Star and Strauss, 1999). Researchers and developers may find that some work cannot or should not be quantifiable, leading to the development of alternative evaluation standards (e.g., worker satisfaction). However, even in occupations where there has been decades of research and development, as is the case with AIED technologies in K-12 education, the vast majority of evaluations for AI deployments have focused on understanding impacts on other stakeholders (e.g., students), without consideration for frontline workers (e.g., teachers’) experiences and perspectives (Holmes, 2023).
In contrast to current workplace measurement approaches, future research could explore the design of new approaches with workers that help make visible the (currently invisible) impacts of AI deployments on their labor. In these efforts, ensuring worker ownership over data collection and disclosure decisions is critical to ensuring that such measurements can actually enable workers to advocate for better technologies and working conditions (or the abandonment of technologies), rather than further surveil (Ming et al., 2023).
@@ 7.4 The Case of Feminized Labor
Redirecting toward a pro-worker future for AI may involve some unique considerations in the context of feminized labor. For instance, many feminized occupations center on emotional and care-oriented labor that may be difficult or impossible to quantify, increasing risks of misquantifications in AI development. Moreover, workers in service-oriented positions (i.e., most feminized positions) are often trained to decenter themselves, becoming a “non-person” in their workplace and to employers (Forsythe, 1993; Ming et al., 2023). Even if AI developers manage to form genuine design partnerships with workers, they may not—without explicit encouragement—advocate for designing technologies that meet their own needs. In past efforts to design technologies with feminized workers (e.g., home health aides, K-12 teachers), these workers often centered clients’ needs and de-centered their own needs by default during design engagements, unless nudged by designers to do so (Holstein et al., 2019; Bartle et al., 2022). We anticipate that, with time, redirecting this trajectory will only increase in importance for AI in feminized occupations. When one considers that many of the most devalued occupations are dominated by women, the majority of whom are women of color, it becomes necessary to recognize the presence and prevalence of AI Failure Loops as a racial, gender, and social justice issue.
@@ 8 Conclusion
Visions for “transformative” workplace AI technologies that undermine workers are not new. Technologists have long envisioned that AI technologies would “fix” staff shortages in feminized occupations. The challenges described in this paper—the development of workplace technology without a grounded understanding of workers and their work—are reminiscent of observations made by Forsythe nearly three decades ago (Forsythe, 1993). Forsythe found that AI developers building expert systems held impoverished understandings of “work,” disregarding social and maintenance activities. Thus, these expert systems failed to model experts’ real-world work practices. Despite the shift from rule-based to machine learning systems, many of Forsythe’s observations persist today.
@@
We would like to thank the reviewers for their excellent reviews that helped strengthen our paper. We are also grateful for the thoughtful feedback shared by Alicia DeVrio, Jodi Forlizzi, Nari Johnson, Joy Ming, Tina Park, and David Widder.